In a Zoo, there are 6 Bengal white tigers and 7 Bengal royal tigers. Out of these tigers, 5 are males and 10 are either Bengal royal tigers or males. Find the number of female Bengal white tigers in the Zoo.


1
Expert's answer
2022-01-05T03:46:41-0500

Solution:

Denotation-

Bengal white tigers = W

Bengal royal tigers = R

Males = M and

Females = F.

Given, "n(W)=6, n(R)=7, n(M)=5, n(F)=8, n(R\cup M)=10"

"\Rightarrow n(T)=n(W)+n(R)=6+7=13."

Also, "n(R\cap M)=n(R)+n(M)-n(R\cup M)"

"\Rightarrow n(R\cap M)=7+5-10=2"

Now, "n(W\cap F)=n(W)+n(F)-n(W\cup F)" ...(i)

Consider "n(W\cup F)"

"=n(T)-n(R\cap M)
\\=13-2
\\=11"

Putting this value in (i).

"n(W\cap F)=6+8-11=3"

Thus, the number of female Bengal white tigers in the Zoo is 3.